# Data-driven controlled subgroup selection in clinical trials

### v1

@@ 1.1 Motivation
Randomised control trials (RCT) are the gold standard for evaluating the safety and efficacy of new treatments. Subgroup discovery and analysis are essential components of the drug development process, as they help identify how different patient groups respond to treatments. This analysis is critical for understanding variations in treatment effects or safety outcomes based on baseline characteristics, and it can lead to personalized treatment plans and improved patient outcomes (lipkovich2017tutorial).
In terms of safety, we would like to uncover subgroups of patients that are less susceptible to adverse events, allowing for the development of safer therapies. One example question that we would like to answer is: “Which participants can be considered low risk and do not require inpatient monitoring after dosing?” This exact question is used in the recent Food and Drug Administration (FDA) guidance document titled “Considerations for the Use of Artificial Intelligence To Support Regulatory Decision-Making” (FDA2025) and it will serve as one of the two motivating examples of our work. In terms of efficacy, subgroup discovery is also critical since it can reveal populations with high treatment effect, paving the way for more personalized and effective treatments. For example, we would like to answer questions such as: “Which subgroups of participants demonstrate more robust improvement in their condition due to the treatment?”
According to alosh2017tutorial, there are three categories of efficacy-based subgroup analysis in drug development:
@@
Much of the current scepticism surrounding the practice of subgroup analysis can be attributed to methodology for exploratory analysis being used incorrectly to try to make inferential claims, for instance by post-hoc specification of the subgroups of interest for which an inferential analysis is conducted (lagakos2006challenge; wang2007Statistics; altman2015subgroup; zhang2015subgroup). As senn1997wisdom put it: “The equivalent in horse-racing would be, rather than betting on a named horse, merely to write ‘first past the post’ on the betting slip and collect the winnings, whatever the result”. However, as we shall see in this paper, recently proposed methodology allows for flexible data-driven selection of subgroups while still giving the strong statistical guarantees necessary for inferential analyses.
@@ 1.2 A new way of approaching the problem of subgroup analysis
To illustrate how this idea differs from more conventional approaches, suppose we have a set of patients (blue dots in Figure 1(a)), and that for each patient we record a response (y-axis), and some baseline biomarker value (x-axis). For the left plot, assume that one has a pre-selected cut-off value for the biomarker (which is often the case in practice); in this example it is chosen as $0.5$. One could then categorize participants as biomarker-negative or biomarker-positive and derive the average group-specific response (represented by horizontal dash-dotted blue lines). This type of subgroup-specific effects is what is considered in many traditional approaches to subgroup analysis; in inferential analyses, their expectations are often the estimands of interest, in exploratory analyses, maximizing the difference between the two groups’ average responses is a common goal, and in supportive analyses, these average effects are what is typically reported (alosh2017tutorial).
@@
To formalize the statistical framework for this problem, consider a regression setting, with $X$ denoting the generic covariate vector and $Y$ the response. Writing $\eta(x):=\mathbb{E}(Y\mid X=x)$, $x\in\mathbb{R}^{d}$, for the population regression function, we are interested in identifying the superlevel-set $S_{\tau}:=\{x\in\mathbb{R}^{d}:\eta(x)\geq\tau\}$ for a pre-specified threshold $\tau\in\mathbb{R}$. More specifically, we are considering the class of settings where the practitioner specifies a nominal limit $\alpha\in(0,1)$ on the Type I error rate and has access to a collection of independent and identically distributed covariate-response pairs $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ following the same distribution as $(X,Y)$. Based on these observations, we seek to select a subset of $\mathbb{R}^{d}$ that is entirely contained within $S_{\tau}$ with probability at least $1-\alpha$. Equivalently, we may be interested in finding a subset of $\mathbb{R}^{d}$ that fully covers $S_{\tau}$ with probability at least $1-\alpha$, or we might seek to address these two tasks simultaneously. This framework may be extended from the pure regression setting (where a single outcome per patient is of interest) to the setting of treatment effects (i.e. where the difference of outcomes across two interventions is of interest), where patients are characterized by triples $(X,T,Y)$, with $T\in\{0,1\}$ a binary control-vs-treatment indicator (hernan2025causal). Details are presented in Section 3.3.
@@ 1.3 Related literature
First and foremost, subgroup-specific effects are historically assessed by the average effect on a subgroup, rather than by the worst case effect on it, as we are doing here through the definition of $S_{\tau}$. This latter definition turns subgroup selection into the problem of conducting inference on the superlevel set of a regression function that has recently been addressed in the context of generalised linear models (GLMs) by wan2024confidence, in smoothness-constrained nonparametric models by reeve2023optimal, and in monotonicity-constrained nonparametric settings by mueller2024isotonic. For the specific case in which the regression function gives a measure of the treatment effect, this framework was also considered in a linear model setting by ballarini2018subgroup and in a univariate, monotonicity-constrained setting by stallard2024testing. In a Bayesian context, this setting of targeting the superlevel set of a linear regression function giving the treatment effect has been considered by schnell2016bayesian, for instance.
@@
Beyond this strengthened guarantee, our framework can be distinguished from much of the existing literature through guaranteeing strong Type I error rate control of the selected set without overly restricting the candidate subsets to be investigated, reflecting recent trends towards principled methods of personalized medicine (lipkovich2024modern). In particular, in response to the dangers of post-hoc selection of subgroups, common guidance on subgroup analysis includes the pre-specification of a small number of biologically plausible subsets based on domain knowledge and correcting any reported results for multiplicity (senn1997wisdom; rothwell2005subgroup; lagakos2006challenge; bretz2014multiplicity). On the other side, the popularity of recently developed data mining tools for detecting subgroups of potential interest illustrate a desire to identify subgroups in a data-driven manner. Some noteworthy tools for such subgroup discovery include: STEPP (bonetti2000graphical; bonetti2004patterns; yip2016subpopulation), a nonparametric smoothing technique to graphically assess treatment effect heterogeneity driven by a single feature which does not directly provide candidate subgroups but can be used to inform the choice thereof; and SIDES (lipkovich2011subgroup) (and its high-dimensional extension SIDEScreen (lipkovich2014strategies)), an iterative method for finding subgroups promising to exhibit large treatment effects with various ways of controlling their complexity. Moreover, a range of tree-based search techniques exist that seek to identify subgroups with differential treatment effects. For instance, su2008interaction; su2009subgroup developed the technique of interaction trees by building on the popular CART-algorithm (breiman1984cart) and using interaction effects between treatment indicator and the covariates as the splitting criterion. A shortcoming of these methods is their tendency to split along covariates with many distinct values, which is improved upon by GUIDE (loh2002regression; loh2015regression). In a similar vein, seibold2016modelbased propose model-based recursive partitioning, a very general method in which the splits are based on independence tests between the partitioning variables and the score function of a parametric model assumed for the effect of covariates and treatment on the primary endpoint. Another important tree-based method is given by causal trees (athey2016recursive) (as well as its random forest version causal forests (wager2018estimation)), which aims to give more accurate estimates of the subgroup-specific effect and use principled splitting strategies (for instance by being designed to reliably select subsets with differential treatment effects and on which the variance when estimating the treatment effect will be small in the case of causal trees). The idea of these methods is to discover parts of the covariate domain that may be driving treatment effect heterogeneity. To confirm these subgroups while controlling the family-wise error rate, one would then need to use an independent fold of the data at a second stage. However, the concern with such splitting strategies is that they reduce power, as fewer observations are available at the confirmation stage, and failure to confirm candidates at this second step leads to no selection whatsoever. Recently, there has been growing interest in methods that replace such rigid two-stage procedures by multi-stage approaches allowing for flexible involvement of the user in the decision at each step (duan2024interactive; cheng2025chiseling).
@@
Moreover, many recent developments for modelling the treatment effect have been made in the area of meta-learners (see lipkovich2024modern for an overview of recent developments), where the idea is to learn the treatment-specific outcome functions and use them to predict the unobservable treatment effect of each patient. This approach has been pioneered by foster2011subgroup who dubbed this technique the “Virtual Twins” method. We investigate how these techniques can be combined with our methods through simulations and real-world applications.
@@ 1.4 Our contribution
reeve2023optimal proved that for these questions and when faced with continuous biomarkers, ensuring control of the Type I error rate across all possible distributions will inevitably lead to trivial power (in the sense that an essentially empty set is selected with high probability) under a large class of data-generating mechanisms. As a result, some assumptions are necessary to yield non-trivial solutions to this problem. To this end, in this paper we present and extend two recently developed tools that build upon two different types of modelling assumptions: one based on generalised linear modelling (GLM) (wan2024confidence) and another on nonparametric monotonicity constraints, which we refer to as Isotonic Subgroup Selection (ISS) (mueller2024isotonic). These methods are applicable in the regression setting, where one is given covariate-response pairs of the form $(X,Y)$, and can be extended to heterogeneous treatment effect settings, i.e. where one has covariate-treatment-response triples $(X,T,Y)$, through inverse propensity weighting. We extend this approach further by building upon a doubly robust way to approximate individual treatment effects (kennedy2023optimal). Furthermore, we carry out an extensive simulation study, considering scenarios designed to closely resemble real clinical trial data as encountered in practice, including situations with both continuous and categorical covariates. The purpose of this analysis is two-fold: to assess the power of the methods, and to evaluate the extent to which they retain Type I error rate control under violations of their respective assumptions.
@@ 2.2 Case Study 2 (efficacy): Subgroup selection for desired treatment response
The second case study focuses on a series of Phase III clinical trials evaluating a pharmaceutical compound targeting psoriatic arthritis (PsA), a chronic inflammatory condition affecting joints, entheses, and skin that can lead to impaired physical function and reduced quality of life (future2). Psoriatic arthritis severity is assessed using a range of clinical measures that capture different aspects of disease activity. Among these are composite endpoints, such as the binary ACR50 score, developed by the American College of Rheumatology (felson1993american). Complementing these are continuous endpoints like tender joint count, swollen joint count, patient global assessment, and physician global assessment, which provide a more granular evaluation of changes within specific disease domains. While our previous case study (Section 2.1) focused on a binary outcome, in this one we aim to demonstrate the generality of the proposed methods by focusing on a continuous endpoint.
Cosentyx (secukinumab) is approved for treating adult patients with active psoriatic arthritis and has been evaluated in several clinical trials, i.e. FUTURE 1-5 (future1; future2; future3; future4; future5). Several studies have analyzed the FUTURE trials to investigate treatment effect heterogeneity with respect to the primary endpoint, ACR50 (Sechidis2021; bornkamp2023predicting; Cardner2023; sechidis2025usingindividualizedtreatmenteffects). With our work, we take a step further by focusing on the identification of patients who experience a treatment effect above a predefined threshold, and we formally pose the following research question.
@@ 3 Methodology
for a nominal Type I error rate $\alpha\in(0,1)$, under suitable assumptions. Note that these properties mean that, with high probability, $\hat{L}_{\alpha}$ (resp. $\hat{L}_{\alpha}^{2\mathrm{S}}$) is a smaller set than $S_{\tau}$ and $\hat{U}_{\alpha}$ (resp. $\hat{U}_{\alpha}^{2\mathrm{S}}$) is a larger set, so that they give lower and upper bounds on $S_{\tau}$ respectively, which is captured by our chosen notation. The superscript $\mathrm{2S}$ refers to the property of having a two-sided bound. These confidence statements are what constitute our strong notion of Type I error rate guarantee. The inferential guarantees considered here even allow for downstream applications under covariate shift. For instance, if a patient has covariate values $X_{0}$ and we find that $X_{0}\in\hat{L}_{\alpha}$, then we may conclude $X_{0}\in S_{\tau}$ and hence $\eta(X_{0})\geq\tau$ (as long as we have not committed a Type I error), no matter whether $X_{0}$ has been drawn from the same distribution as $X_{1},\ldots,X_{n}$ or has been picked in any other way. Corresponding conclusions can be made when instead $\hat{U}_{\alpha}$, $\hat{L}_{\alpha}^{2\mathrm{S}}$ or $\hat{U}_{\alpha}^{2\mathrm{S}}$ are considered. This is in stark contrast to the conclusions offered by conducting inference on the average effect on a subset (as compared in Figure 1), where one can at best draw conclusions for the average value $\eta(X_{0})$ but also that only when $X_{0}$ is sampled from the same distribution as $X_{1},\ldots,X_{n}$. In other words, in our framework, the covariate values in the training data $X_{1},\ldots,X_{n}$ do not need to be representative for the target population to which the patient with values $X_{0}$ belongs.

### v2

@@ 1.1 Motivation
Randomised control trials (RCT) are the gold standard for evaluating the safety and efficacy of new treatments. Subgroup discovery and analysis are essential components of the drug development process, as they help identify how different patient groups respond to treatments. This analysis is critical for understanding variations in treatment effects or safety outcomes based on baseline characteristics, and it can lead to personalized treatment plans and improved patient outcomes (Lipkovich et al., 2017).
In terms of safety, we would like to uncover subgroups of patients that are less susceptible to adverse events, allowing for the development of safer therapies. One example question that we would like to answer is: “Which participants can be considered low risk and do not require inpatient monitoring after dosing?” This exact question is used in the recent Food and Drug Administration (FDA) guidance document titled “Considerations for the Use of Artificial Intelligence To Support Regulatory Decision-Making” (U.S. Food and Drug Administration, 2025) and it will serve as one of the two motivating examples of our work. In terms of efficacy, subgroup discovery is also critical since it can reveal populations with high treatment effect, paving the way for more personalized and effective treatments. For example, we would like to answer questions such as: “Which subgroups of participants demonstrate more robust improvement in their condition due to the treatment?”
According to Alosh et al. (2017), there are three categories of efficacy-based subgroup analysis in drug development:
@@
Much of the current scepticism surrounding the practice of subgroup analysis can be attributed to methodology for exploratory analysis being used incorrectly to try to make inferential claims, for instance by post-hoc specification of the subgroups of interest for which an inferential analysis is conducted (Lagakos, 2006; Wang et al., 2007; Altman, 2015; Zhang et al., 2015). As Senn and Harrell (1997) put it: “The equivalent in horse-racing would be, rather than betting on a named horse, merely to write ‘first past the post’ on the betting slip and collect the winnings, whatever the result”. However, as we shall see in this paper, recently proposed methodology allows for flexible data-driven selection of subgroups while still giving the strong statistical guarantees necessary for inferential analyses.
@@ 1.2 A new way of approaching the problem of subgroup analysis
To illustrate how this idea differs from more conventional approaches, suppose we have a set of patients (blue dots in Figure 1(a)), and that for each patient we record a response (y-axis), and some baseline biomarker value (x-axis). For the left plot, assume that one has a pre-selected cut-off value for the biomarker (which is often the case in practice); in this example it is chosen as $0.5$. One could then categorize participants as biomarker-negative or biomarker-positive and derive the average group-specific response (represented by horizontal dash-dotted blue lines). This type of subgroup-specific effects is what is considered in many traditional approaches to subgroup analysis; in inferential analyses, their expectations are often the estimands of interest, in exploratory analyses, maximizing the difference between the two groups’ average responses is a common goal, and in supportive analyses, these average effects are what is typically reported (Alosh et al., 2017).
@@
To formalize the statistical framework for this problem, consider a regression setting, with $X$ denoting the generic covariate vector and $Y$ the response. Writing $\eta(x):=\mathbb{E}(Y\mid X=x)$, $x\in\mathbb{R}^{d}$, for the population regression function, we are interested in identifying the superlevel-set $S_{\tau}:=\{x\in\mathbb{R}^{d}:\eta(x)\geq\tau\}$ for a pre-specified threshold $\tau\in\mathbb{R}$. More specifically, we are considering the class of settings where the practitioner specifies a nominal limit $\alpha\in(0,1)$ on the Type I error rate and has access to a collection of independent and identically distributed covariate-response pairs $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ following the same distribution as $(X,Y)$. Based on these observations, we seek to select a subset of $\mathbb{R}^{d}$ that is entirely contained within $S_{\tau}$ with probability at least $1-\alpha$. Equivalently, we may be interested in finding a subset of $\mathbb{R}^{d}$ that fully covers $S_{\tau}$ with probability at least $1-\alpha$, or we might seek to address these two tasks simultaneously. This framework may be extended from the pure regression setting (where a single outcome per patient is of interest) to the setting of treatment effects (i.e. where the difference of outcomes across two interventions is of interest), where patients are characterized by triples $(X,T,Y)$, with $T\in\{0,1\}$ a binary control-vs-treatment indicator (Hernan and Robins, 2025). Details are presented in Section 3.3.
@@ 1.3 Related literature
First and foremost, subgroup-specific effects are historically assessed by the average effect on a subgroup, rather than by the worst case effect on it, as we are doing here through the definition of $S_{\tau}$. This latter definition turns subgroup selection into the problem of conducting inference on the superlevel set of a regression function that has recently been addressed in the context of generalised linear models (GLMs) by Wan et al. (2024), in smoothness-constrained nonparametric models by Reeve et al. (2023), and in monotonicity-constrained nonparametric settings by Müller et al. (2025). For the specific case in which the regression function gives a measure of the treatment effect, this framework was also considered in a linear model setting by Ballarini et al. (2018) and in a univariate, monotonicity-constrained setting by Stallard (2024). In a Bayesian context, this setting of targeting the superlevel set of a linear regression function giving the treatment effect has been considered by Schnell et al. (2016), for instance.
@@
Beyond this strengthened guarantee, our framework can be distinguished from much of the existing literature through guaranteeing strong Type I error rate control of the selected set without overly restricting the candidate subsets to be investigated, reflecting recent trends towards principled methods of personalized medicine (Lipkovich et al., 2024). In particular, in response to the dangers of post-hoc selection of subgroups, common guidance on subgroup analysis includes the pre-specification of a small number of biologically plausible subsets based on domain knowledge and correcting any reported results for multiplicity (Senn and Harrell, 1997; Rothwell, 2005; Lagakos, 2006; Bretz and Westfall, 2014). On the other side, the popularity of recently developed data mining tools for detecting subgroups of potential interest illustrate a desire to identify subgroups in a data-driven manner. Some noteworthy tools for such subgroup discovery include: STEPP (Bonetti and Gelber, 2000, 2004; Yip et al., 2016), a nonparametric smoothing technique to graphically assess treatment effect heterogeneity driven by a single feature which does not directly provide candidate subgroups but can be used to inform the choice thereof; and SIDES (Lipkovich et al., 2011) (and its high-dimensional extension SIDEScreen (Lipkovich and Dmitrienko, 2014)), an iterative method for finding subgroups promising to exhibit large treatment effects with various ways of controlling their complexity. Moreover, a range of tree-based search techniques exist that seek to identify subgroups with differential treatment effects. For instance, Su et al. (2008, 2009) developed the technique of interaction trees by building on the popular CART-algorithm (Breiman et al., 1984) and using interaction effects between treatment indicator and the covariates as the splitting criterion. A shortcoming of these methods is their tendency to split along covariates with many distinct values, which is improved upon by GUIDE (Loh, 2002; Loh et al., 2015). In a similar vein, Seibold et al. (2016) propose model-based recursive partitioning, a very general method in which the splits are based on independence tests between the partitioning variables and the score function of a parametric model assumed for the effect of covariates and treatment on the primary endpoint. Another important tree-based method is given by causal trees (Athey and Imbens, 2016) (as well as its random forest version causal forests (Wager and Athey, 2018)), which aims to give more accurate estimates of the subgroup-specific effect and use principled splitting strategies (for instance by being designed to reliably select subsets with differential treatment effects and on which the variance when estimating the treatment effect will be small in the case of causal trees). The idea of these methods is to discover parts of the covariate domain that may be driving treatment effect heterogeneity. To confirm these subgroups while controlling the family-wise error rate, one would then need to use an independent fold of the data at a second stage. However, the concern with such splitting strategies is that they reduce power, as fewer observations are available at the confirmation stage, and failure to confirm candidates at this second step leads to no selection whatsoever. Recently, there has been growing interest in methods that replace such rigid two-stage procedures by multi-stage approaches allowing for flexible involvement of the user in the decision at each step (Duan et al., 2024; Cheng et al., 2025).
@@
Moreover, many recent developments for modelling the treatment effect have been made in the area of meta-learners (see Lipkovich et al. (2024) for an overview of recent developments), where the idea is to learn the treatment-specific outcome functions and use them to predict the unobservable treatment effect of each patient. This approach has been pioneered by Foster et al. (2011) who dubbed this technique the “Virtual Twins” method. We investigate how these techniques can be combined with our methods through simulations and real-world applications.
@@ 1.4 Our contribution
Reeve et al. (2023, Prop. 1) proved that for these questions and when faced with continuous biomarkers, ensuring control of the Type I error rate across all possible distributions will inevitably lead to trivial power (in the sense that an essentially empty set is selected with high probability) under a large class of data-generating mechanisms. As a result, some assumptions are necessary to yield non-trivial solutions to this problem. To this end, in this paper we present and extend two recently developed tools that build upon two different types of modelling assumptions: one based on generalised linear modelling (GLM) (Wan et al., 2024) and another on nonparametric monotonicity constraints, which we refer to as Isotonic Subgroup Selection (ISS) (Müller et al., 2025). These methods are applicable in the regression setting, where one is given covariate-response pairs of the form $(X,Y)$, and can be extended to heterogeneous treatment effect settings, i.e. where one has covariate-treatment-response triples $(X,T,Y)$, through inverse propensity weighting. We extend this approach further by building upon a doubly robust way to approximate individual treatment effects (Kennedy, 2023). Furthermore, we carry out an extensive simulation study, considering scenarios designed to closely resemble real clinical trial data as encountered in practice, including situations with both continuous and categorical covariates. The purpose of this analysis is two-fold: to assess the power of the methods, and to evaluate the extent to which they retain Type I error rate control under violations of their respective assumptions.
@@ 2.2 Case Study 2 (efficacy): Subgroup selection for desired treatment response
The second case study focuses on a series of Phase III clinical trials evaluating a pharmaceutical compound targeting psoriatic arthritis (PsA), a chronic inflammatory condition affecting joints, entheses, and skin that can lead to impaired physical function and reduced quality of life (McInnes et al., 2015). Psoriatic arthritis severity is assessed using a range of clinical measures that capture different aspects of disease activity. Among these are composite endpoints, such as the binary ACR50 score, developed by the American College of Rheumatology (Felson et al., 1993). Complementing these are continuous endpoints like tender joint count, swollen joint count, patient global assessment, and physician global assessment, which provide a more granular evaluation of changes within specific disease domains. While our previous case study (Section 2.1) focused on a binary outcome, in this one we aim to demonstrate the generality of the proposed methods by focusing on a continuous endpoint.
Cosentyx (secukinumab) is approved for treating adult patients with active psoriatic arthritis and has been evaluated in several clinical trials, i.e. FUTURE 1-5 (Mease et al., 2015; McInnes et al., 2015; Nash et al., 2018; Kivitz et al., 2019; Mease et al., 2018). Several studies have analyzed the FUTURE trials to investigate treatment effect heterogeneity with respect to the primary endpoint, ACR50 (Sechidis et al., 2021; Bornkamp et al., 2024; Cardner et al., 2023; Sechidis et al., 2025b). With our work, we take a step further by focusing on the identification of patients who experience a treatment effect above a predefined threshold, and we formally pose the following research question.
@@ 3 Methodology
for a nominal Type I error rate $\alpha\in(0,1)$, under suitable assumptions. Note that these properties mean that, with high probability, $\hat{L}_{\alpha}$ (resp. $\hat{L}_{\alpha}^{2\mathrm{S}}$) is a smaller set than $S_{\tau}$ and $\hat{U}_{\alpha}$ (resp. $\hat{U}_{\alpha}^{2\mathrm{S}}$) is a larger set, so that they give lower and upper bounds on $S_{\tau}$ respectively, which is captured by our chosen notation. The superscript $\mathrm{2S}$ refers to the property of having a two-sided bound. These confidence statements are what constitute our strong notion of Type I error rate guarantee.
The inferential guarantees considered here even allow for downstream applications under covariate shift. For instance, if $\hat{L}_{\alpha}\subseteq S_{\tau}$ (i.e. we have not committed a Type I error) and we have a patient with covariate values $X_{0}$ such that $X_{0}\in\hat{L}_{\alpha}$, then it follows that $X_{0}\in S_{\tau}$ and hence $\eta(X_{0})\geq\tau$, no matter whether $X_{0}$ has been drawn from the same distribution as $X_{1},\ldots,X_{n}$ or has been picked in any other way. Corresponding conclusions can be made when instead $\hat{U}_{\alpha}$, $\hat{L}_{\alpha}^{2\mathrm{S}}$ or $\hat{U}_{\alpha}^{2\mathrm{S}}$ are considered. This is in stark contrast to the conclusions offered by conducting inference on the expected effect averaged over a subset (as compared in Figure 1), where one can at best draw conclusions for the average value of $\eta(X_{0})$ but also that only when $X_{0}$ is sampled from the same distribution as $X_{1},\ldots,X_{n}$. In other words, in our framework, the covariate values in the training data $X_{1},\ldots,X_{n}$ do not need to be representative for the target population to which the patient with values $X_{0}$ belongs. This is particularly important when considering that exclusion criteria imposed as part of the recruitment for clinical trials may cause the distribution of the trial population to deviate from that of the target population. Whenever such exclusion criteria only affect the covariate distribution, our statistical guarantees remain intact, a robustness typically not exhibited by conventional approaches.